OpenStudy (anonymous):

For the following number, give the opposite, the reciprocal, and the absolute value. -4/5

OpenStudy (compassionate):

\[\frac{ 4 }{ 5 } and \frac{ 5 }{ 4 } and \left| \frac{ 4 }{ 5 } \right|\] The opposite is just positive, the reciprocal is flipped; making it positive, the absolute value is the spaces from zero, you can't be negative spaces from zero, thus, it is positive.